### v1

Abstract
We consider statistical inference problems under uncertain equality constraints, and provide asymptotically valid uncertainty estimates for inferred parameters. The proposed approach leverages the implicit function theorem and primal-dual optimality conditions for a particular problem class. The motivating application is multi-dimensional raking, where observations are adjusted to match marginals; for example, adjusting estimated deaths across race, county, and cause in order to match state all-race all-cause totals. We review raking from a convex optimization perspective, providing explicit primal-dual formulations, algorithms, and optimality conditions for a wide array of raking applications, which are then leveraged to obtain the uncertainty estimates. Empirical results show that the approach obtains, at the cost of a single solve, nearly the same uncertainty estimates as computationally intensive Monte Carlo techniques that pass thousands of observed and of marginal “draws” through the entire raking process.
keywords:
raking, delta method, implicit function theorem
{MSCcodes}
46N10, 62D05, 62E20
1 Introduction
Classic asymptotic uncertainty relies on properties of estimating functions. Given $n$ observations $y_{i}$ that depend on $\beta\in\mathbb{R}^{p}$, an estimating function  [19] is a set of $p$ equations satisfying
(1) $G\left(\beta\right)=\sum_{i=1}^{n}G\left(\theta;y_{i}\right)=0\quad\mbox{where}\quad E\left[G\left(\beta\right)\right]=0\quad\forall\theta.$
Asymptotic uncertainty uses (1) along with standard regularity conditions  [17]. The canonical choice of $G$ is the gradient of the negative log likelihood for the data:
(2) $\min_{\beta}f\left(\beta;y\right)\quad\Rightarrow\quad G\left(\beta;y\right)=\nabla_{\beta}f\left(\beta;y\right).$
We now consider a generalized system of estimating equations in higher dimensional space:
(3) $F\left(\beta,\lambda;y,s\right)=0,\quad F:\mathbb{R}^{p+k}\rightarrow\mathbb{R}^{p+k}$
where $\lambda$ is another parameter set, while $s$ is another random input, similar to $y$. This extension is motivated by equality constraints with random right hand sides
(4) $\min_{\beta}f\left(\beta;y\right)\quad\mbox{s.t.}\quad A\beta=s.$
The classic Lagrangian approach then yields a system of form  (3).
Problem  (4) is of fundamental importance in raking, which is widely used in survey sciences and epidemiology. Given a set of noisy observations $y$, and marginal constraints $s$, the goal is to find nearby $\beta$ that minimizes a fitness functional, such as entropic distance, subject to the constraints. Both the detailed observations $y$ and aggregate observations $s$ are subject to uncertainty, which should be rigorously captured by uncertainty quantification methods for $\beta$. The main contribution of this paper is to develop an efficient and rigorous way to propagate uncertainty for the entire class (3).
Convex Preliminaries
Define
$\overline{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}.$
For a function $f:\mathbb{R}^{p}\rightarrow\overline{\mathbb{R}}$, we define
$\text{epi}\left(f\right)=\{\left(\beta,\alpha\right):f\left(\beta\right)\leq\alpha\}\subset\mathbb{R}^{p+1}.$
A function $f:\mathbb{R}^{p}\rightarrow\overline{\mathbb{R}}$ is
  • closed when $\text{epi}\left(f\right)$ is a closed set
  • proper when $f\not\equiv\infty$ and does not take the value of $\infty$- ∞ anywhere
  • convex when $\text{epi}\left(f\right)$ is a convex set.
For any function $f:\mathbb{R}^{p}\rightarrow\overline{\mathbb{R}}$, the convex conjugate is defined by
(5) $f^{*}\left(z\right)=\sup_{\beta}z^{T}\beta-f\left(\beta\right).$
For closed convex $f$, we have $f^{**}=f$. When $f$ and $f^{*}$ are closed and differentiable, Fenchel’s equality gives us a relation between the gradient maps:
$\nabla_{z}f^{*}\left(\cdot\right)=\nabla_{\beta}f\left(\cdot\right)^{-1}.$
Conjugacy implies simple calculus rules that we use later for specific examples:
  • Scaling:
    (6) $\left(\alpha f\right)^{*}\left(z\right)=\sup_{\beta}z^{T}\beta-\alpha f\left(\beta\right)=\alpha f^{*}\left(z/\alpha\right).$
  • Coordinate sums:
    (7) $f\left(x\right)=\sum_{i}f_{i}\left(x_{i}\right)\quad\Rightarrow\quad f^{*}\left(z\right)=\sum_{i}f_{i}^{*}\left(z_{i}\right).$
  • Scaled coordinate sums:
    (8) $f_{\alpha}\left(x\right)=\sum_{i}\alpha_{i}f_{i}\left(x_{i}\right)\quad\Rightarrow\quad f_{\alpha}^{*}\left(z\right)=\sum_{i}\alpha_{i}f_{i}^{*}\left(z_{i}/\alpha_{i}\right).$
Given a convex program
(9) $\min_{\beta}f\left(\beta\right)\quad\text{s.t.}\quad A\beta=s,$
the canonical Lagrangian is given by
(10) $\mathcal{L}\left(\beta,\lambda\right)=f\left(\beta\right)+\lambda^{T}\left(A\beta-s\right).$
The original problem is equivalent to solving
$\min_{\beta}\sup_{\lambda}\mathcal{L}\left(\beta,\lambda\right)$
while the dual is obtained by interchanging the order of optimization
(11) $\sup_{\lambda}\min_{\beta}f\left(\beta\right)+\lambda^{T}\left(A\beta-s\right)$ $=\sup_{\lambda}-\lambda^{T}s-\sup_{\beta}-\left(A^{T}\lambda\right)^{T}\beta-f\left(\beta\right)$
$=\sup_{\lambda}-\lambda^{T}s-f^{*}\left(-A^{T}\lambda\right).$
When $f$ is also differentiable, basic technical conditions (see e.g. [13, Theorem 11.39]) allow the primal and dual solution to be simultaneously obtained by solving a system of nonlinear equations characterizing the stationarity of the Lagrangian (10):
(12) $0\quad=\quad F\left(\beta,\lambda;y,s\right)=\begin{bmatrix}\nabla_{\beta}f\left(\beta\right)+A^{T}\lambda\\ A\beta-s\end{bmatrix}$
When the dual solution $\lambda^{*}$ is available,  (12) can be used to recover the primal
(13) $\beta^{*}=\left(\nabla_{\beta}f\right)^{-1}\left(-A^{T}\lambda^{*}\right).$
Roadmap
The paper is organized as follows. The description and some analysis of the raking problem is in Section 2, while the main result on uncertainty propagation is presented in Section 3. Section 4 provides a simple illustration of the approach on a small synthetic dataset. We take a deeper dive into a 3-dimensional raking application for mortality estimation in Section 5. Conclusions follow in Section 6.
2 Raking formulations and methods
Raking is widely used in survey inference to adjust the sampling weights of the cases in a contingency table in order to match the marginal totals of the adjusted weights to the corresponding totals for the population  [16]. Let us suppose that we have a population $\mathcal{U}=\left\{1,\cdots,i,\cdots,p\right\}$ and are interested in estimating the population total of a variable $Y$ which takes the values $y_{i}$ for $i=1,\cdots,p$
(14) $\mathcal{Y}=\sum_{i\in\mathcal{U}}y_{i}.$
While we typically cannot access $y_{i}$ for all $i\in\mathcal{U}$, we can use values taken by individuals from a sample $S\subseteq\mathcal{U}$. If we denote $p\left(.\right)$ a sampling design, that is a probability measure on the set of all possible samples $S\subseteq\mathcal{U}$, the inclusion probability is defined by $\pi_{i}=\sum_{S:i\in S}p\left(S\right)$ and we denote $d_{i}=\frac{1}{\pi_{i}}$. The classic estimator for $\mathcal{Y}$, also called the Horvitz–Thompson–Narain estimator  [12, 11], is then
(15) $\hat{\mathcal{Y}}=\sum_{i\in S}\frac{y_{i}}{\pi_{i}}=\sum_{i\in S}d_{i}y_{i}.$
We may also have access to information on $k$ auxiliary variables. We define by
(16) $x_{i}=\left(x_{i}^{1},\cdots,x_{i}^{j},\cdots,x_{i}^{k}\right)^{T}\in\mathbb{R}^{k}$
the vector of the values taken by the $k$ auxiliary variables for an individual $i$. The values taken by $x_{i}^{j}$ for individual $i$ and auxiliary variable $j$ are not known, but we suppose that the margin vector
(17) $\sum_{i\in\mathcal{U}}x_{i}=\left(x^{1},\cdots,x^{j},\cdots,x^{k}\right)^{T}\in\mathbb{R}^{k}$
is known. Our objective is to replace the weights $d_{i}$ in  (15) by raking weights $w_{i}$ such that
(18) $\sum_{i\in S}w_{i}x_{i}=\sum_{i\in\mathcal{U}}x_{i}.$
An estimator for $\mathcal{Y}$ is then
(19) $\hat{\mathcal{Y}}_{raking}=\sum_{i\in S}w_{i}y_{i}.$
In the classic raking problem, the variable of interest is $\mathcal{Y}$ and we are interested in computing an estimate of $\mathcal{Y}$ with associated uncertainty. In this paper, the value of $\mathcal{Y}$ is known and our interest is in the weights $w_{i}$ and their associated uncertainty.
2.1 One-dimensional raking
We are interested in the case where the sample $S$ is equal to the population $\mathcal{U}$ and the population total $\mathcal{Y}$ is known, but the values taken by the $y_{i}$ must be adjusted such that their sum is equal to the known total. The weights $d_{i}$ are then equal to $1$. There is just one auxiliary variable $x^{1}$, equal to the data vector $y$, and the single marginal total $\sum_{i\in\mathcal{U}}x_{i}^{1}$ is equal to the known sum of the observations. Instead of looking for the weights $w_{i}$, we are interested in directly finding the raked values $\beta_{i}=w_{i}y_{i}$. The scalar raking problem is given by
(20) $\min_{\beta}f\left(\beta;y\right)\quad\text{s.t.}\quad 1^{T}\beta=s.$
From (11), the dual is given by
(21) $\max_{\lambda}-\lambda s-f^{*}(-\lambda 1).$
When $f$ is the entropic distance function, from simple calculus and (7) we have
$f(\beta;y)=\sum_{i=1}^{p}\beta_{i}\log\left(\frac{\beta_{i}}{y_{i}}\right)-\beta_{i}+y_{i},\qquad f^{*}(z)=y^{T}\left(e^{z}-1\right)\text{ and }\nabla_{z}f^{*}\left(z\right)=y\odot e^{z}.$
where $\odot$ denotes the element-wise product between length-$p$ vectors $y$ and $e^{z}$. The dual problem now has a closed form solution
$-s+1^{T}\nabla_{z}f^{*}\left(-\lambda 1\right)=0\quad\Rightarrow\quad\lambda^{*}=-\log\left(\frac{s}{1^{T}y}\right)$
and from (13) we have
(22) $\beta^{*}=\left[\nabla_{\beta}f\left(\beta,y\right)\right]^{-1}\left(-\lambda^{*}1\right)\quad\Rightarrow\quad\log\left(\beta^{*}/y\right)=-\lambda^{*}1\quad\Rightarrow\quad\beta^{*}=y\frac{s}{1^{T}y}.$
Adjusting all the estimates using the ratio between the constraint and the sum of the observations solves the problem. In higher-dimensional settings, the closed-form solution is no longer available, but the idea of solving the dual problem proves very useful, since for common distance measures $f$ it is smooth, unconstrained, and has smaller dimension.
2.2 Two-dimensional raking and iterative proportional fitting (IPF) algorithm
In the two-dimensional raking problem, we are given linear combinations of the data (e.g. $n$ row sums and $m$ column sums) as constraints. We then have $k=m+n$ auxiliary variables. The raking problem now takes the form
(23) $\min_{\beta}f(\beta;y)\quad\text{s.t.}\quad\beta^{T}1=s_{1}\text{ and }\beta 1=s_{2},$
where $\beta,y\in\mathbb{R}^{m\times n}$, $s_{1}\in\mathbb{R}^{n}$ and $s_{2}\in\mathbb{R}^{m}$. An algorithm to solve (23) with $f$ the entropic distance function
(24) $f\left(\beta;y\right)=\sum_{i=1}^{m}\sum_{j=1}^{n}\beta_{i,j}\log\left(\frac{\beta_{i,j}}{y_{i,j}}\right)-\beta_{i,j}+y_{i,j}$
was first proposed by [4] and further developed by [15], though these early works focused on the iterative raking procedure rather than the underlying formulation (23)-(24). The calibration method was further developed by [6] and [7] to adjust samples on known population totals. A review on the method and optimization formulations is found in [5], with the exception of the explicit dual and primal-dual relationships developed here.
The Lagrangian takes the form
(25) $\min_{\beta}\max_{\lambda_{1},\lambda_{2}}\mathcal{L}\left(\beta,\lambda_{1},\lambda_{2}\right):=f\left(\beta;y\right)+\lambda_{1}^{T}\left(\beta^{T}1-s_{1}\right)+\lambda_{2}^{T}\left(\beta 1-s_{2}\right),$
with the dual problem given by
(26) $\sup_{\lambda_{1},\lambda_{2}}-\lambda_{1}^{T}s_{1}-\lambda_{2}^{T}s_{2}-f^{*}(-\lambda_{1}1^{T}-1\lambda_{2}^{T})$
where by (7) we have
$f^{*}(z)=\mbox{trace}\left(Y^{T}\left(\exp\left(Z\right)-\mathbf{1}\right)\right),$
with $\mathbf{1}$ a matrix of ones of the same dimension as $Z$. In particular, given the dual solution $\lambda^{*}$, the primal solution is obtained from (13):
(27) $\beta_{i,j}^{*}=y_{i,j}e^{-\lambda_{1,j}^{*}-\lambda_{2,i}^{*}}.$
Each raked value $\beta_{i,j}$ is thus equal to the corresponding observation $y_{i,j}$ multiplied by a factor that depends only on the row $i$ and a factor that depends only on the column $j$. This view leads to the classic Iterative Proportional Fitting algorithm (IPF,  [4, 15]), which corresponds to alternating minimization for the dual variables $\lambda_{1}$ and $\lambda_{2}$, tracked in the primal space using (27). Specifically, IPF is initialized with
(28) $\beta_{i,j}^{\left(0\right)}=y_{i,j}$
and then alternates between column and row updates
(29) $\beta_{i,j}^{\left(2t-1\right)}=\beta_{i,j}^{\left(2t-2\right)}\frac{s_{1,j}}{\sum_{i=1}^{m}\beta_{i,j}^{\left(2t-2\right)}},\qquad\beta_{i,j}^{\left(2t\right)}=\beta_{i,j}^{\left(2t-1\right)}\frac{s_{2,i}}{\sum_{j=1}^{n}\beta_{i,j}^{\left(2t-1\right)}}.$
The IPF algorithm is a particular way to exploit the primal-dual relationship (23)-(26) and correspondence (27) , but it breaks down when we change the optimization problem (23) by modifying the objective $f$, including adding weights, or incorporating additional terms. We use algorithms that attack the dual directly in more general formulations.
2.3 Multi-dimensional weighted raking
Here, we study more in detail the general case with $k$ auxiliary variables which are a linear combination of the data. We suppose that we have a vector of detailed observations $y\in\mathbb{R}^{p}$. The margin vector $\sum_{i\in\mathcal{U}}x_{i}$ is denoted $s$ and the constraints become $A\beta=s$. We introduce a vector of weights $q\in\mathbb{R}^{p}$ that do not depend on $y$ or $s$ and denote $\frac{1}{q}=\left(\frac{1}{q_{1}},\cdots,\frac{1}{q_{i}},\cdots,\frac{1}{q_{p}}\right)^{T}$ the vector containing the inverses of the weights. Denoting
(30) $f_{1/q}\left(\beta;y\right)=\sum_{i=1}^{p}\frac{1}{q_{i}}f_{i}\left(\beta_{i},y_{i}\right)$
we have the primal problem
(31) $\beta^{*}=\operatorname*{arg\,min}_{\beta}f_{1/q}\left(\beta;y\right)\quad\mbox{s.t.}\quad A\beta=s$
and corresponding dual
(32) $\min_{\lambda}\lambda^{T}s+f_{1/q}^{*}\left(-A^{T}\lambda\right)$
with $f_{1/q}^{*}$ defined as in (8), yielding a simpler expression for the dual:
(33) $\min_{\lambda}\lambda^{T}s+\sum_{i}\frac{1}{q_{i}}f_{i}^{*}\left(-q_{i}a_{i}^{T}\lambda\right).$
Taking the derivative with respect to $\lambda$, we get the dual optimality conditions:
(34) $s=A\nabla_{z}f^{*}\left(-q\odot A^{T}\lambda\right).$
Once we have the optimal $\lambda^{*}$, (34), we can recover the primal solution. From the Lagrangian, we have
(35) $0=\nabla_{\beta}f_{1/q}(\beta)+A^{T}\lambda$
and so given $\lambda^{*}$ we have
(36) $\beta^{*}=\nabla_{z}f_{1/q}^{*}(-A^{T}\lambda^{*})=\nabla_{z}f^{*}(-q\odot A^{T}\lambda^{*}).$
Solving the dual
To solve (34) we use Newton’s method. Denote
(37) $\Phi\left(\lambda\right)=Ay-A\nabla_{z}f^{*}\left(-q\odot A^{T}\lambda;y\right)\text{ and }\hat{s}=Ay.$
Then the non-linear system can be written
(38) $\Phi\left(\lambda\right)-\hat{s}+s=0.$
We start at $\lambda^{0}=0$ and at each iteration $t$, we solve the system
(39) $\nabla_{\lambda}\Phi\left(\lambda^{t}\right)\left(\lambda^{t}-\lambda^{t+1}\right)=\Phi\left(\lambda^{t}\right)-\hat{s}+s$
with
(40) $\nabla_{\lambda}\Phi\left(\lambda^{t}\right)=A\left[\text{diag}\left(q\odot\nabla^{2}_{\lambda,z}f^{*}\left(-q\odot A^{T}\lambda^{t};y\right)\right)\right]A^{T}.$
The Jacobian matrix $\nabla_{\lambda}\Phi\left(\lambda^{t}\right)$ may not be invertible. For example, in the two-dimensional case, if we translate the two-dimensional table $\beta_{i,j}$ into a vector such that $\beta_{i,j}=\beta_{\left(j-1\right)m+i}$, the two constraints $\beta^{T}1=s_{1}\text{ and }\beta 1=s_{2}$ become $A\beta=s$ where $A$ is an $\left(n+m\right)\times\left(mn\right)$ matrix with rank $n+m-1$. However, as the two margins $s_{1}$ and $s_{2}$ need to have equal sums (that is $1^{T}s_{1}=1^{T}s_{2}$) for the raking problem to have a solution, the last constraint $\beta_{m.}^{T}1=s_{2m}$ is a linear combination of the $n+m-1$ other constraints and can be dropped from the constraint matrix $A$. We then get an $\left(n+m-1\right)\times\left(mn\right)$ matrix $A$ which has full rank and the corresponding Jacobian matrix $\nabla_{\lambda}\Phi\left(\lambda^{t}\right)$ is a symmetric invertible matrix. At each iteration, we thus have
(41) $\lambda^{t+1}=\lambda^{t}-\left[\nabla_{\lambda}\Phi\left(\lambda^{t}\right)\right]^{-1}\left(\Phi\left(\lambda^{t}\right)-\hat{s}+s\right).$
Provided that the rows of the constraint matrix $A$ form a basis, the Jacobian matrix $\nabla_{\lambda}\Phi\left(\lambda^{t}\right)$ will always be invertible. For the classic raking problem, the question of the existence of solutions was previously addressed in  [18, 5]. In the application presented in Section 5, the size of matrix $A$ is at most $2289\times 6096$, leading to a Jacobian matrix of size $2289\times 2289$. We solve the linear system using the sparse conjugate gradient method implemented in the scipy.sparse.linalg.cg function. Another possible issue is that the iteration $t$ may lead to a solution $\lambda^{t+1}$ such that $-q\odot A^{T}\lambda^{t+1}$ is not in the domain of $\nabla^{2}_{\lambda,z}f^{*}$. In that case, an interior-point method may be used  [1, 5].
To implement the raking method, we must choose the distance function $f\left(\beta;y\right)$. Until now, we implicitly supposed that for a fixed value $y$, the function $f\left(.;y\right):\mathbb{R}^{p}\rightarrow\mathbb{R}^{+}$ is positive, differentiable and strictly convex such that $f\left(y;y\right)=0$. A simple choice is the $\mathcal{L}_{2}$ norm
(42) $f\left(\beta;y\right)=\sum_{i=1}^{p}\frac{1}{2}\left(\beta_{i}-y_{i}\right)^{2}.$
In this case, the solution of the dual problem is obtained by solving a single linear system for $\lambda$:
(43) $s=A\left[y-q\odot A^{T}\lambda,\right]$
and the raked values are given by
(44) $\beta^{*}=y-q\odot A^{T}\lambda^{*}.$
From the formula, it is possible to obtain negative raked values even if all observations are positive, which is a non-starter when the variable of interest must stay positive, for example when raking death counts or death rates.
A family of distance functions is proposed by  [6] and  [7], who, in our notation, start with the gradient of the conjugate:
(45) $\nabla_{z}f^{*}(z)\left(z;y\right)=y\odot\left(1+\alpha z\right)^{\frac{1}{\alpha}}\;\Rightarrow\;f\left(\beta;y\right)=\sum_{i=1}^{p}\frac{1}{\alpha}\left[\frac{y_{i}}{\alpha+1}\left(\frac{\beta_{i}}{y_{i}}\right)^{\alpha+1}-\beta_{i}+c\left(y_{i}\right)\right]$
where $c\left(y_{i}\right)$ is a function that may depend on the observations $y_{i}$ but not on the unknown raked values $\beta_{i}$. In the case of $\alpha=1$, we get the $\chi^{2}$ distance
(46) $f\left(\beta;y\right)=\sum_{i=1}^{p}\frac{1}{2y_{i}}\left(\beta_{i}-y_{i}\right)^{2}.$
The solution of the dual problem is obtained by solving the linear system:
(47) $s=A\left[y\odot\left(1-q\odot A^{T}\lambda\right)\right]$
and the raked values are equal to
(48) $\beta^{*}=y\odot\left(1-q\odot A^{T}\lambda^{*}\right).$
Taking $\alpha=0$, we recover the familiar entropic distance
(49) $\nabla_{z}f^{*}(z)\left(z;y\right)=y\odot e^{z}\quad\Rightarrow\quad f\left(\beta;y\right)=\sum_{i=1}^{p}\beta_{i}\log\left(\frac{\beta_{i}}{y_{i}}\right)-\beta_{i}+y_{i}.$
The solution of the dual problem is obtained by iteratively solving the non-linear system:
(50) $s=A\left[y\odot\exp\left(-q\odot A^{T}\lambda\right)\right]$
and the raked values are given by
(51) $\beta^{*}=y\odot\exp\left(-q\odot A^{T}\lambda^{*}\right)$
One particular advantage of this last distance function is that the raked values will always remain of the same sign as the corresponding observations.
Finally, if we want the raked values to stay bounded (e.g. $l_{i}\leq\beta_{i}\leq h_{i}$), we can use the logit distance
(52) $f\left(\beta;y\right)=\sum_{i=1}^{p}\left(\beta_{i}-l_{i}\right)\log\frac{\beta_{i}-l_{i}}{y_{i}-l_{i}}+\left(h_{i}-\beta_{i}\right)\log\frac{h_{i}-\beta_{i}}{h_{i}-y_{i}}.$
All of these distance functions are convex and differentiable, with convex and differentiable conjugates, so the method developed in this section applies. The choice of the distance function is further discussed in  [3].
3 Uncertainty
In the classic raking problem, a quantity of interest is the variance of the raking estimator  (19). The quantities $y_{i},i\in S$ are known and the sources of uncertainty will be how the sample $S$ is selected from the population $\mathcal{U}$ and the uncertainty on the marginal totals $\sum_{i\in\mathcal{U}}x_{i}$. The variance of the estimator  (15) proposed by  [12, 11] is
(53) $\mathbb{V}\left(\hat{\mathcal{Y}}\right)=\sum_{k,l\in\mathcal{U}}\left(\pi_{kl}-\pi_{k}\pi_{l}\right)\frac{y_{k}}{\pi_{k}}\frac{y_{l}}{\pi_{l}}.$
A review on methods to use priors on the margins to compute the posterior distribution for the total $\mathcal{Y}$ is given in  [14]. However, in this paper, the total sum $\mathcal{Y}$ and its variance are known and we are interested in using the known variances and covariances on the observations $y$ and the margins $s$ to estimate the unknown covariance matrix of the raked values $\beta$.
In the case where $f$ is equal to the $\mathcal{L}_{2}$ norm, the solution of the system  (43) is
(54) $\beta^{*}=y-q\odot A^{T}\lambda^{*}\text{ with }A\left[\operatorname{diag}\left(q\right)\right]A^{T}\lambda^{*}=Ay-s.$
Denoting $\Phi=A\left[\text{diag}\left(q\right)\right]A^{T}$ and $\Phi^{-1}$ its inverse, we have
(55) $\lambda^{*}=\Phi^{-1}\left(Ay-s\right)\text{ and }\beta^{*}=\left(I-\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}A\right)y+\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}s.$
We denote $\Sigma_{y}$ the $p\times p$ covariance matrix of the observation vector $y$, $\Sigma_{s}$, the $k\times k$ covariance matrix of the aggregate observations vector $s$ and $\Sigma_{ys}$ the $p\times k$ covariance matrix of $y$ and $s$. The covariance matrix of the estimated raked values is then
(56) $\Sigma_{\beta^{*}}$ $=\left(I-\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}A\right)\Sigma_{y}\left(I-\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}A\right)^{T}$
$+\left(\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}\right)\Sigma_{s}\left(\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}\right)^{T}$
$+2\left(I-\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}A\right)\Sigma_{ys}\left(\left[\text{diag}\left(q\right)\right]A^{T}\Phi^{-1}\right)^{T}.$
In the general case, the raked values cannot be written as a linear combination of the observations and we leverage the implicit function theorem to get a method for uncertainty propagation. We go back to the Lagrangian optimality conditions:
(57) $F\left(\beta,\lambda;y,s\right)=0$
where
(58) $F\left(\beta,\lambda;y,s\right)=\begin{bmatrix}\frac{1}{q}\odot\nabla_{\beta}f\left(\beta;y\right)+A^{T}\lambda\\ A\beta-s.\end{bmatrix}$
The solution of this system can be written as
(59) $\beta^{*}=\phi\left(y;s\right)\text{ with }\phi:\mathbb{R}^{p+k}\rightarrow\mathbb{R}^{p}.$
Theorem 3.1 (Covariance of the raked values $\beta^{*}$).
Let $\theta$ denote the expectation of the observations and margins, $\Sigma$ denote their covariance matrix, and $\phi^{\prime}_{\theta}$ denote the matrix of the partial derivatives of $\phi$ with respect to the observations $y$ and margins $s$ taken at $\theta$. Then the covariance matrix of the raked values $\beta^{*}$ is given by
$\Sigma_{\beta}=\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}.$
Proof 3.2.
Suppose that we have $n$ vectors of observations $\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$ taking values in $\mathbb{R}^{p+k}$. We can compute their sample mean $\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}$ and through the central limit theorem, we have
(60) $\sqrt{n}\left(\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}-\theta\right)\rightarrow\mathcal{N}\left(0,\Sigma\right)$
where $\theta=\mathbb{E}\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$, $\Sigma$ is the covariance matrix
(61) $\Sigma=\begin{pmatrix}\Sigma_{y}&\Sigma_{ys}\\ \Sigma_{ys}&\Sigma_{s}\end{pmatrix}$
and $\mathcal{N}\left(0,\Sigma\right)=T$ denotes the multivariate normal distribution with mean 00 and covariance matrix $\Sigma$. If the function $\phi$ mapping the detailed observations and the aggregate observations to the raked estimates is differentiable at $\theta$, then we have
(62) $\sqrt{n}\left(\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{n}\end{matrix}\right)-\phi\left(\theta\right)\right)\rightarrow\phi^{\prime}_{\theta}\left(T\right)$
where
(63) $\phi^{\prime}_{\theta}\left(T\right)=\begin{pmatrix}\frac{\partial\phi_{1}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{1}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial s_{k}}\left(\theta\right)\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial\phi_{p}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial s_{k}}\left(\theta\right)\\ \frac{\partial\phi_{p+1}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+1}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p+1}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+1}}{\partial s_{k}}\left(\theta\right)\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial\phi_{p+k}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+k}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p+k}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+k}}{\partial s_{k}}\left(\theta\right)\end{pmatrix}T,$
and so $\sqrt{n}\left(\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{n}\end{matrix}\right)-\phi\left(\theta\right)\right)$ converges weakly to a multivariate normal distribution with mean 00 and covariance $\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}$. To relate the general result to the specific example (56), with the $\chi^{2}$ distance, (57) is
(64) $\begin{pmatrix}I&\left[\text{diag}\left(q\odot y\right)\right]A^{T}\\ A&0\end{pmatrix}\begin{pmatrix}\beta\\ \lambda\end{pmatrix}=\begin{pmatrix}y\\ s\end{pmatrix}.$
If $\beta_{1}^{*}$ solves $F\left(\beta,\lambda;y_{1},s_{1}\right)=0$ and $\beta_{2}^{*}$ solves $F\left(\beta,\lambda;y_{2},s_{2}\right)=0$, then $\beta_{1}^{*}+\alpha\beta_{2}^{*}$ solves
$F\left(\beta,\lambda;y_{1}+y_{2},s_{1}+\alpha s_{2}\right)=0.$
From this, we can conclude that
(65) $\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{s}\end{matrix}\right)=\frac{1}{n}\sum_{i=0}^{n}\phi\left(\begin{matrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{matrix}\right).$
From  (59), the covariance matrix of $\phi\left(\begin{matrix}y^{(i)}\\ s^{(i)}\end{matrix}\right)$ is equal to the covariance matrix of the raked values $\Sigma_{\beta}$. Thus we have through the central limit theorem
(66) $\sqrt{n}\left(\frac{1}{n}\sum_{i=0}^{n}\phi\left(\begin{matrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{matrix}\right)-\phi\left(\theta\right)\right)\rightarrow\mathcal{N}\left(0,\Sigma_{\beta}\right).$
Gathering  (62) and  (66), we finally get
(67) $\Sigma_{\beta}=\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}.$
To compute $\Sigma_{\beta}$ in the general case, we now need to compute the partial derivatives $\frac{\partial\phi_{j}}{\partial y_{i}}$ and $\frac{\partial\phi_{j}}{\partial s_{i}}$. We use the Implicit Function Theorem [10], adapted below to our notation.
Theorem 3.3 (Implicit Function Theorem).
Assume that $S$ is an open subset of $\mathbb{R}^{2p+2k}$ and that $F:S\rightarrow\mathbb{R}^{p+k}$ is a function of class $C^{1}$. Assume also that $\left(\beta^{*},\lambda^{*};y_{0},s_{0}\right)$ is a point in $S$ such that
(68) $F\left(\beta^{*},\lambda^{*};y_{0},s_{0}\right)=0,\quad\det D_{\beta,\lambda}F\left(\beta^{*},\lambda^{*};y_{0},s_{0}\right)\neq 0$
where $D_{\beta,\lambda}F$ is the matrix of partial derivatives
(69) $D_{\beta,\lambda}F=\begin{pmatrix}\frac{\partial F_{1}}{\partial\beta_{1}}&\ldots&\frac{\partial F_{1}}{\partial\beta_{p}}&\frac{\partial F_{1}}{\partial\lambda_{1}}&\ldots&\frac{\partial F_{1}}{\partial\lambda_{k}}\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial F_{p+k}}{\partial\beta_{1}}&\ldots&\frac{\partial F_{p+k}}{\partial\beta_{p}}&\frac{\partial F_{p+k}}{\partial\lambda_{1}}&\ldots&\frac{\partial F_{p+k}}{\partial\lambda_{k}}\end{pmatrix}.$
Then we have the following:
  1. i.
    There exist $r_{0},r_{1}>0$ such that for every $\left(y,s\right)\in\mathbb{R}^{p+k}$ with
    $\left\lVert\left(y,s\right)-\left(y_{0},s_{0}\right)\right\rVert<r_{0},$
    there is a unique $\left(\beta,\lambda\right)\in\mathbb{R}^{p+k}$ such that
    (70) $\left\lVert\left(\beta,\lambda\right)-\left(\beta^{*},\lambda^{*}\right)\right\rVert<r_{1}\text{ and }F\left(\beta,\lambda;y,s\right)=0.$
    Thus (70) implicitly defines a function $\left(\beta,\lambda\right)=\phi\left(y,s\right)$ for $\left(y,s\right)\in\mathbb{R}^{p+k}$ near $\left(y_{0},s_{0}\right)$, with $\left(\beta,\lambda\right)=\phi\left(y,s\right)$ close to $\left(\beta^{*},\lambda^{*}\right)$. In particular $\phi\left(y_{0},s_{0}\right)=\left(\beta^{*},\lambda^{*}\right)$.
  2. ii.
    Moreover, the function $\phi:B\left(r_{0};y,s\right)\rightarrow B\left(r_{1};\beta,\lambda\right)\subset\mathbb{R}^{p+k}$ from part (i.) above is of class $C^{1}$, and its derivatives may be determined by differentiating the identity
    (71) $F\left(y,s;\phi\left(y,s\right)\right)=0$
    (a consequence of the definition of $\phi$) and solving to find the partial derivatives of $\phi$.
When differentiating $F\left(y,s;\phi\left(y,s\right)\right)=0$ at the solution $\left(\beta^{*},\lambda^{*}\right)$, we get
(72) $\left[D_{\beta,\lambda}F\left(y,s;\beta^{*},\lambda^{*}\right)\right]\left[D_{y,s}\phi\left(y,s\right)\right]=-\left[D_{y,s}F\left(y,s;\beta^{*},\lambda^{*}\right)\right].$
The unknown matrix that we need to compute to get the gradient is
(73) $D_{y,s}\phi=\begin{pmatrix}\frac{\partial\beta}{\partial y}&\frac{\partial\beta}{\partial s}\\ \frac{\partial\lambda}{\partial y}&\frac{\partial\lambda}{\partial s}\end{pmatrix}$
and we can compute easily the two other matrices. Denoting $F_{1}=\frac{1}{q}\odot\nabla_{\beta}f\left(\beta;y\right)+A^{T}\lambda$ and $F_{2}=A\beta-s$, we have
(74) $D_{\beta,\lambda}F=\begin{pmatrix}\frac{\partial F_{1}}{\partial\beta}&\frac{\partial F_{1}}{\partial\lambda}\\ \frac{\partial F_{2}}{\partial\beta}&\frac{\partial F_{2}}{\partial\lambda}\end{pmatrix}=\begin{pmatrix}\left[\text{diag}\left(\frac{1}{q}\right)\right]\nabla^{2}_{\beta}f\left(\beta;y\right)&A^{T}\\ A&0_{k\times k}\end{pmatrix}$
and
(75) $D_{y,s}F=\begin{pmatrix}\frac{\partial F_{1}}{\partial y}&\frac{\partial F_{1}}{\partial s}\\ \frac{\partial F_{2}}{\partial y}&\frac{\partial F_{2}}{\partial s}\end{pmatrix}=\begin{pmatrix}\left[\text{diag}\left(\frac{1}{q}\right)\right]\nabla^{2}_{\beta y}f\left(\beta;y\right)&0_{p\times k}\\ 0_{k\times p}&-I_{k\times k}\end{pmatrix}.$
Due to the choice of the distance function $f$, the Hessian matrices $\nabla^{2}_{\beta}f\left(\beta;y\right)$ and $\nabla^{2}_{\beta y}f\left(\beta;y\right)$ are diagonal. If we prune the constraints to ensure $A$ has full rank as described above, both $D_{\beta,\lambda}F$ and $D_{y,s}F$ are invertible. The system  (72) has then a unique solution.
4 Simulated Example
We apply the raking procedure to a small synthetic 2-dimensional example with two categorical variables $X_{1}$ and $X_{2}$. We suppose that $X_{1}$ can take $m=3$ values $\left[1;2;3\right]$ and that $X_{2}$ can take $n=5$ values $\left[1;2;3;4;5\right]$ and we generate $m\times n$ values following a uniform distribution over the interval $\left[2;3\right]$ to fill the initial matrix $\beta_{0}$. We compute the margins $s_{1}=\beta_{0}^{T}1$ and $s_{2}=\beta_{0}1$. Then we add some random noise to get the observations $y_{0}=\beta_{0}+\mathcal{N}\left(\mu=0,\sigma=0.1\right)$. We then generate $N=1000$ samples following a multivariate normal distribution with mean $y_{0}$ and covariance $\Sigma$. The off-diagonal elements of $\Sigma$ are equal to $0.01$ while the diagonal elements of $\Sigma$ are equal to $\Sigma_{k,k}=0.1\times k$ for $k=1,\cdots,m\times n$. In this particular example, we suppose that there is no uncertainty on the margins. We can then compute the mean $\bar{y}$ of the observations $y^{(i)}$ and apply the raking procedure to the mean to get $\beta^{*}=\phi\left(\bar{y},s_{1},s_{2}\right)$. The results are shown in Figure 1. The blue dots represent the initial observations while the orange squares represent the raked values. The uncertainty on the observations and the raked values is marked by a segment of length two times the standard deviation and centered on the observations or the raked values.
Figure 1: Initial observations (blue dots) and corresponding raked values (orange squares) with their corresponding uncertainties. The raked values are close to the initial observations, but have lower uncertainties as there is no uncertainty on the margins.
We also show how much the initial value of observation $\bar{y}_{3,4}$ affects the corresponding raked values $\beta^{*}_{i,j},i=1,\cdots,m;j=1,\cdots n$ in Figure 2. The raked value affected most corresponds to the initial observation $\beta^{*}_{3,4}$. An increase in $\bar{y}_{3,4}$ is associated with an increase of $\beta^{*}_{3,4}$. Conversely, increasing $\bar{y}_{3,4}$ will decrease the raked values on the same row or column $\beta^{*}_{3,j},j=1,2,3,5$ and $\beta^{*}_{i,4},i=1,2$.
Figure 2: Influence of one observation $\bar{y}_{3,4}$ on all the raked values. As expected, the corresponding raked value $\beta^{*}_{3,4}$ is the most affected while the correlations are negative for the raked values $\beta^{*}_{3,j}$ and the $\beta^{*}_{i,4}$.
Similarly, we show how much the raked value $\beta^{*}_{3,4}$ depends on the initial observations $\bar{y}_{i,j},i=1,\cdots,m;j=1,\cdots,n$ in Figure 3. The observation with the bigger influence is the one that corresponds to the raked value $\bar{y}_{3,4}$. An increase in $\bar{y}_{3,4}$ is associated with an increase of $\beta^{*}_{3,4}$. Conversely, an increase of the initial values on the same row or column $\bar{y}_{3,j},j=1,2,3,5$ and $\bar{y}_{i,4},i=1,2$ will decrease the value of $\beta^{*}_{3,4}$.
Figure 3: Influence of all the observations on one raked value $\beta^{*}_{3,4}$. As expected, the corresponding observation $\bar{y}_{3,4}$ has the biggest influence while the correlations are negative for the observations $\bar{y}_{3,j}$ and $\bar{y}_{i,4}$.
Finally, we compare in Figure 4 the raked value $\beta^{*}$ obtained when raking the mean of the $1000$ samples and when raking each of the samples and computing the mean of the raked values. We also compare the variance of $\beta^{*}$ computed with the uncertainty propagation method described in section 3 with the variance computed by raking the $1000$ samples of the observations and computing the sample variance of the corresponding raked values (that is, using an empirical process to quantify uncertainty). We get very similar results using both methods, showing that the asymptotic method leveraging the implicit function theorem is very effective.
Figure 4: Mean (left) and variance (right) of the raked values computed with the delta method or with all the samples. The method we developed gives nearly identical results as the Monte Carlo method.
5 Application
Our dataset [9] contains for each US state, year, sex, and age group the estimated number of deaths by cause, race and ethnicity and county. These estimates are raked to aggregate estimates, typically all-race state-wide cause-specific and all-cause counts, while trying to preserve key relational information in the detailed estimates, as described in detail below.
There are $I=3$ different causes (communicable, maternal, neonatal, and nutritional diseases (Comm.); non-communicable diseases (NCD); and injuries (Inj.)) and $J=5$ different races (White; Black; American Indian or Alaska Native (AIAN); Asian or Pacific Islander (API); and Hispanic/Latino (Hisp.)) in the dataset. We denote $y_{i,j,k}$ the number of deaths for cause $i=1,\cdots,I$, race $j=1,\cdots,J$ and county $k=1,\cdots,K$. We denote $i=0$ for the sum of deaths over all causes, $j=0$ for the sum of deaths over all races, and $k=0$ for the sum of deaths over all counties, that is the number of deaths at the state level. The only aggregate data known are provided by the GBD  [2] and are the total number of deaths for the state $s_{0}=y_{0,0,0}$ and the total number of deaths due to cause $i$ for the state $s_{i}=y_{i,0,0}$.
The modeling challenge here is to also ensure internal consistency, guaranteeing that:
  1. 1.
    For each county $k$, the sum of all-races deaths by cause $i$ is equal to the all-causes, all-races deaths: $\sum_{i=1}^{I}\beta_{i,0,k}=\beta_{0,0,k}$.
  2. 2.
    For each county $k$, the sum of all-causes deaths by race $j$ is equal to to the all-causes, all-races deaths: $\sum_{j=1}^{J}\beta_{0,j,k}=\beta_{0,0,k}$.
  3. 3.
    For each race $j$ and each county $k$, the sum of deaths by cause is equal to the all-causes deaths: $\sum_{i=1}^{I}\beta_{i,j,k}=\beta_{0,j,k}$.
  4. 4.
    For each cause $i$ and each county $k$, the sum of deaths by race is equal to the all-races deaths: $\sum_{j=1}^{J}\beta_{i,j,k}=\beta_{i,0,k}$.
In a previous study [8], the raking was not done on race and ethnicity; one-dimensional raking was done using the closed form solution  (22) and IPF was used for the two-dimensional raking. A follow-up study [9] used race and ethnicity but without the GBD margins. To extend [8] and [9] to the case where raking is done for the three variables cause, race and ethnicity, and county, the workflow described in Algorithm 1 was first proposed. It used one- and two-dimensional raking to achieve the requirements described above.
Algorithm 1 Initial procedure: 4-steps workflow
  Minimize $\sum_{k=1}^{K}f\left(\beta_{0,0,k},y_{0,0,k}\right)$ under the constraint $\sum_{k=1}^{K}\beta_{0,0,k}=s_{0}$.
  Minimize $\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)$ under the constraints $\sum_{k=1}^{K}\beta_{i,0,k}=s_{i}$ for all $i=1,\cdots,I$ and $\sum_{i=1}^{I}\beta_{i,0,k}=\beta_{0,0,k}$ for all $k=1,\cdots,K$.
  for $k=1,\cdots,K$ do
     Minimize $\sum_{j=1}^{J}f\left(\beta_{0,j,k},y_{0,j,k}\right)$ under the constraint $\sum_{j=1}^{J}\beta_{0,j,k}=\beta_{0,0,k}$.
  end for
  for $k=1,\cdots,K$ do
     Minimize $\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$ under the constraints $\sum_{i=1}^{I}\beta_{i,j,k}=\beta_{0,j,k}$ for all $j=1,\cdots,J$ and $\sum_{j=1}^{J}\beta_{i,j,k}=\beta_{i,0,k}$ for all $i=1,\cdots,I$.
  end for
Using the formulations and methods in the paper, we can rake on cause, race and county in one step, using any distance function we like, using weighting as necessary, and obtain uncertainty through a single solve, avoiding re-solving the problem 1000 times (once for each draw). The corresponding workflow is described in Algorithm 2.
Algorithm 2 Proposed procedure: 1-step workflow
  Minimize $\sum_{k=1}^{K}f\left(\beta_{0,0,k},y_{0,0,k}\right)+\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)+\sum_{k=1}^{K}\sum_{j=1}^{J}f\left(\beta_{0,j,k},y_{0,j,k}\right)+\sum_{k=1}^{K}\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$ under the constraints:
  • $\sum_{k=1}^{K}\beta_{0,0,k}=s_{0}$;
  • $\sum_{k=1}^{K}\beta_{i,0,k}=s_{i}\text{ for }i=1,\cdots,I$;
  • $\sum_{i=1}^{I}\beta_{i,0,k}-\beta_{0,0,k}=0\text{ for }k=1,\cdots,K$;
  • $\sum_{j=1}^{J}\beta_{0,j,k}-\beta_{0,0,k}=0\text{ for }k=1,\cdots,K$;
  • $\sum_{i=1}^{I}\beta_{i,j,k}-\beta_{0,j,k}=0\text{ for }j=1,\cdots,J\text{ and }k=1,\cdots,K$;
  • $\sum_{j=1}^{J}\beta_{i,j,k}-\beta_{i,0,k}=0\text{ for }i=1,\cdots,I\text{ and }k=1,\cdots,K$.
In this proposed workflow, if we denote $l=(I+1)(J+1)(k-1)+(I+1)j+i+1$ and we take $s_{l}=0$ for $l>i$, we can rewrite the problem as
(76) $\min_{\beta}\sum_{l=1}^{p}\frac{1}{q_{l}}f\left(\beta_{l};y_{l}\right)\quad\text{s.t.}\quad A\beta=s$
where $p=(I+1)(J+1)K$ and $A$ is a $k\times p$ sparse matrix containing only 00s, $1$s and $-1$s. In the initial formulation of Algorithm 2, there are $1+I+\left(I+J+2\right)K$ constraints. However, the first constraint $\sum_{k=1}^{K}\beta_{0,0,k}=s_{0}$ and the last $K$ constraints $\sum_{j=1}^{J}\beta_{I,j,k}-\beta_{I,0,k}=0$ are a linear combination of the other constraints. We can drop them and finally get a $k\times p$ matrix $A$ that has full rank, with $k=I+\left(I+J+1\right)K$. All the linear systems in Sections 2 and 3 will then have a unique solution.
In Figure 5, we compute the raking procedure for Delaware, both sexes, and age group 25-to-30-year-old, using the entropic distance and equal weights. We compare the reported values of the death rates (number of deaths divided by the population for each race and county) with the corresponding raked values for both workflows for all 72 cases (3 counties, 5 races and all races, 3 causes and all causes). The 1-step approach gives raked values slightly closer to the initial values than the 4-steps workflow, meaning that the raked estimates stay closer to the original reported values while satisfying the requirements. This improvement is easier to see for larger reported death rates on the linear scale plot in Figure 5.
Figure 5: Initial rate of death and corresponding raked values for the two proposed workflows. The raked values obtained with the proposed 1-step workflow are usually closer to the initial observations than the raked values obtained with the previous 4-step workflow.
We now compare the raked values with their associated uncertainty to the initial values. To better compare the different causes, races and counties on the same scale, we compare mortality rates. In Figure 6, Figure 7 and Figure 8, we look at the initial values and the raked values for Delaware, both sexes and age group 0-to-1-year-old. The uncertainty is represented by a segment of length two standard deviations and centered on the estimated value. Races $3$ and $4$ with much smaller population numbers have the largest uncertainties for both the initial and the raked values. This uncertainty on the raked values is significantly bigger than the uncertainty on the initial values. The initial uncertainty on the margins $s$ (that is on the GBD values) are much larger than the uncertainties on the observations $y$, and as a result the raking process significantly increases the final uncertainty on these raked values.
Figure 6: Initial and raked values with associated uncertainty ordered by cause. The raked values are close to the initial observations, but have much bigger uncertainties, as the uncertainty on the margins is much bigger than the uncertainty reported for the observations.
Figure 7: Initial and raked values with associated uncertainty ordered by race. The raked values are close to the initial observations, but have much bigger uncertainties as the uncertainty on the margins is much bigger than the uncertainty reported for the observations.
Figure 8: Initial and raked values with associated uncertainty ordered by county. The raked values are close to the initial observations, but have much bigger uncertainties as the uncertainty on the margins is much bigger than the uncertainty reported for the observations.
To estimate how the initial values influence the raked values, we look at the initial value for cause “injuries” (Inj), race group Black and New Castle County (denoted $y_{2,2,2}$). Then we compute the corresponding values of the gradient $\frac{\partial\beta_{i,j,k}}{\partial y_{2,2,2}}$ for $i=0,\cdots,3$, $j=0,\cdots,5$ and $k=1,2,3$. The corresponding values of the gradient are shown in Figure 9. The most affected raked value is the one with the same cause, race and county as the initial value considered. All-causes deaths for race group Black and New Castle County and all-races deaths for cause “injuries” and New Castle County will be positively influenced by variations in the initial value $y_{2,2,2}$ while other causes and races will be negatively influenced. The raked values in the other counties will not be much affected by changes in the value of $y_{2,2,2}$.
Figure 9: Influence of the initial value of deaths for cause injuries, race group Black and New Castle County on all the raked values. As expected, the corresponding raked value $\beta^{*}_{2,2,2}$ is the most affected, followed by the raked values for the partial totals $\beta^{*}_{0,2,2}$ and the $\beta^{*}_{2,0,2}$.
Conversely, we want to know what are the observation points that have the most influence on the final estimated raked values. We look at the raked value for cause “injuries”, race group Black and New Castle County (denoted $\beta_{2,2,2}$). Then we compute the corresponding values of the gradient $\frac{\partial\beta_{2,2,2}}{\partial y_{i,j,k}}$ and $\frac{\partial\beta_{2,2,2}}{\partial s_{i}}$ for $i=0,\cdots,3$, $j=0,\cdots,5$ and $k=1,2,3$. The corresponding values of the gradient are shown in Figure 10. The most important observation point is the one with the same cause, race and county as the raked value considered. Increases in all-causes and all-races deaths will increase the raked value, while increases in other causes and races will decrease it. The initial values for the other counties will have a smaller influence on the raked value. For the margins, increases in all-causes, all-races state mortality and in all-races state mortality due to injuries will increase the raked value while increases in all-races state mortality due to other causes will decrease it.
Figure 10: Influence of all the initial values and aggregated initial values on the raked value for cause injuries, race group Black and New Castle County. As expected, the corresponding observation $\bar{y}_{2,2,2}$ is the most influential, followed by the margin for the cause “injuries”.
Finally, we compute the variance of the estimated raked values using two methods: First, we take $100$ samples of the detailed and aggregate observations, apply the raking procedure to each sample, and compute the sample mean and the sample variance of the corresponding raked values. Second, we compute the sample mean of the initial data set, apply the raking procedure to the sample mean, and use the delta method and the implicit function theorem to propagate the variance of the observations to get the variance of the raked values. In Figure 11, we looked at the raked values for Delaware, both sexes, and all age groups and compared the raked values and their standard deviation computed with both methods. We get very similar results with the delta method as we would have with a Monte Carlo technique that rakes all draws, gaining huge savings in compute.
Figure 11: Comparison between the means (left) and the standard deviations (right) of the raked values computed with the delta method or with all the samples. The method we developed gives nearly identical results as the Monte Carlo method that requires solving a raking problem for each draw of reported estimates and marginal totals.
6 Conclusions
In this paper, we have considered the multi-dimensional raking problem, where observations are adjusted to match marginals; for example, adjusting estimated deaths across race, county, and cause to match state all-race all-cause totals. We have provided asymptotically valid uncertainty estimates for inferred parameters that correctly account for uncertain constraints. To capture the uncertainty in both reported data and marginals, we leveraged the Lagrangian framework for optimization with linear constraints, and used the implicit function theorem to obtain the asymptotic uncertainty estimates using primal-dual optimality conditions. The approach obtains nearly the same uncertainty estimates as Monte Carlo estimation techniques that pass thousands of marginal ‘draws’ through the multi-dimensional raking estimation process, at the cost of a single solve.
Acknowledgments
This work was funded by the Bill and Melinda Gates Foundation.